# HyperAIRI: a plug-and-play algorithm for precise hyperspectral image reconstruction in radio interferometry

### v1

@@ 5.4 Reconstruction results
Figure 6: Reconstructed images of the field SB9442-35 at the frequencies 817 MHz (left) and 1069 MHz (right), obtained using the monochromatic algorithms AIRI (top), and uSARA (bottom) (Wilber et al., 2023b, a). Each panel is overlaid with zooms on selected regions: zoom (i) is centred on the “dancing ghosts”; zooms (ii) and (iv) show regions with both extended and compact radio sources; zoom (iii) contains the star-forming galaxy NGC 7090; zooms (v)-(viii) focus on 4 compact sources.
6 Validation on real ASKAP observations
In this section, we revisit the ASKAP measurements of the field SB9442-35 previously imaged by Wilber et al. (2023a, b) with uSARA and AIRI. To evaluate the performance of proposed hyperspectral algorithms, Hyper-uSARA and HyperAIRI, we support our analysis with the reconstructions obtained by their monochromatic counterparts, as well as the hyperspectral CLEAN variant in WSClean.
6.1 Data details for the field SB9442-35
@@
Figure 7: Reconstructed images of the field SB9442-35 at the frequencies 817 MHz (left) and 1069 MHz (right), obtained using the WSClean with joint-channel deconvolution. The zoom regions are identical to those in Fig. 6.
The hyperspectral measurements used in this experiment are part of the ASKAP EMU Pilot Survey (Norris et al., 2021). In this survey, each Scheduling Block (SB) provides 36 different measurement sets, each corresponding to a different primary beam. For the real data experiment, we select the measurement set of Beam 35 of SB9442, centred at the radio source PKS 2130-538 of the galaxy cluster Abell 3785, known as the “dancing ghosts”. The observation spans 10 hours covering the frequency band 800-1088 MHz with a frequency step size of 1 MHz, yielding 288 spectral channels. The selected hyperspectral measurements were calibrated with the ASKAPsoft (Hotan et al., 2021) and are accessible through the CSIRO ASKAP Science Data Archive (Chapman et al., 2015).
6.2 Imaging settings
Following the imaging considerations of Wilber et al. (2023a, b), the measurement set is divided into 8 spectral windows, combining 36 channels each. The visibilities in each spectral window are concatenated together during imaging. The target image cube thus contains 8 spectral channels. Each channel has a pixel size of $4096\times 4096$ and a cell size of 2.20 arcsec, achieving a final FoV of $2.57^{\circ}\times 2.57^{\circ}$, corresponding to a super-resolution factor of 2.22 at the highest-frequency channel and 2.91 at the lowest-frequency channel. The $w$-correction is enabled for all algorithms to account for $w$-effect in wide-field imaging. Briggs weighting with the robustness parameter fixed to 0.25-0.25 is considered.
[…]
To evaluate the spectral index maps from the reconstructed image cubes, we follow the considerations of Wilber et al. (2023a, b). For each image cube, a smoothing Gaussian kernel was applied, with a full width at half maximum (FWHM) of 20 arcsec for WSClean and 5 arcsec for the remaining algorithms. Subsequently, the logarithm of the pixel intensities of these channels are fitted to a linear function with the slope to be the spectral index map, as shown in (23).
@@ 6.3 Reconstruction results
Figure 8: Reconstructed images of the field SB9442-35 at the frequencies 817 MHz (left) and 1069 MHz (right), obtained using the HyperAIRI (top) and Hyper-uSARA (bottom). The zoom regions are identical to those in Fig. 6.
@@
We examine the reconstructions provided by the different algorithms in Fig. 6-8, at both lowest (left panels) and highest (right panels) frequencies. Fig. 6 exhibits AIRI (top panels) and uSARA (bottom panels) images from Wilber et al. (2023b, a). Fig. 7 presents WSClean images. Fig. 8 displays HyperAIRI (top panels) and Hyper-uSARA (bottom panels) images. In each panel, we consider zooms on eight selected regions as follows: region (i) centred at the “dancing ghosts”; regions (ii) and (iv) corresponding to areas containing both extended and compact radio sources; region (iii) centred on the star-forming galaxy NGC 7090; and regions (v)-(viii) centred on compact sources with varying morphologies.
@@
Compared with their monochromatic counterparts, the proposed hyperspectral algorithms recover significantly more faint background sources, capture finer details of the radio emission, and exhibit improved inter-channel consistency. Among the hyperspectral approaches, HyperAIRI achieves distinguished source sharpness and recovers the largest number of compact sources, albeit with some noise-like residual patterns suggesting some extent of data over-fitting. On the other hand, Hyper-uSARA fails to capture some of the faint compact background sources due to its joint sparsity regularisation, enforcing pixel values of sources, detected in some channels only, to be set to zero. WSClean reconstructions are smoother by design, with some faint sources appear to be buried in the noise. While one could further optimise the noise-level scaling factors on a per-channel basis as in Wilber et al. (2023a, b) or consider per-channel visibility calibration to better balance background residuals against source recovery, these processing strategies lie beyond the scope of this experiment.
@@
A detailed examination of the zoom regions highlights substantial differences in the imaging precision of the different algorithms in terms of angular resolution, dynamic range and the ability to capture faint emission, and spectral correlation. Firstly, the superior angular resolution achieved by HyperAIRI at both channels compared to the benchmark is exemplified by zoom regions (vi), (vii) and (viii), with Hyper-uSARA ranking second. Both algorithms reveal inner cores and jets of the selected radio galaxies, whereas their monochromatic counterparts and even WSClean provide a blurred depiction of the same sources. Notably, HyperAIRI resolves the inner core of the radio galaxy in region (vi), and distinctly separates the two compact sources at the centre of region (viii). Secondly, the enhanced sensitivity and dynamic-range of the proposed hyperspectral algorithms is showcased in zoom region (v). Both HyperAIRI and Hyper-uSARA successfully recover the faint jets emanating from the two hotspots. The faint features are completely absent in the monochromatic reconstructions, and are not revealed clearly in WSClean’s images. Thirdly, consistency of spectral behaviour achieved by the proposed hyperspectral algorithms is exemplified by the V-shaped emission in region (ii), the star-forming galaxy NGC 7090 in region (iii), and the two-sided jet emission and nearby point sources in region (iv), with faint structures preserved across channels. Examining region (iii) at the highest frequency, monochromatic algorithms fail to recover NGC 7090’s fusiform morphology. For WSClean, the source is buried in the noise. By contrast, both HyperAIRI and Hyper-uSARA reliably restore NGC 7090’s elongated form, with the former providing a clearer depiction of the diffuse emission.
@@
Figure 9: Residual dirty images zoomed at the “dancing ghosts”, for channel 1 (top row) and 8 (bottom row) from the reconstructions produced by the algorithms shown in Figs. 6-8
@@
Figure 10: The spectral index maps of the “dancing ghosts” evaluated from the reconstructions of different algorithms. Panels (a) and (b) are extracted from the reconstructions of Wilber et al. (2023a, b). The maps are computed after convolving channel images with a Gaussian beam of 20 arcsec for WSClean and 5 arcsec for the other reconstructions.
In what follows, we provide a deeper analysis of the recovery of the “dancing ghosts” in zoom region (i), including data fidelity and the source’s spectral index map. The source is the most prominent in this field, notable for its complex morphology. It comprises the emission of two active galactic nuclei (AGN): the first (WISEA J213406.70-533418.7) at the centre of the north bridge, and the second (WISEA J213417.69-533811.1) in the southeast filament (Norris et al., 2021; Velović et al., 2023). All hyperspectral algorithms provide consistent reconstructions across channels. In particular, two jets from the two lobes of the first AGN respectively pointing towards the northeast exhibit similar shapes in both the first and last channels. Hyper-uSARA offers sharper edges and lower background noise, albeit with minor wavelets artefacts around the jets and lobes of the second AGN. HyperAIRI tends to deliver slightly higher resolution and recover more of the neighbouring compact point-like sources than Hyper-uSARA. However, the PnP algorithm seems to depict the north bridge with a chain of dots, which is likely a reconstruction artefact. This could be indicative of training dataset bias, with half of the images simulated from optical astronomical images dominated by point-like sources. For insight into the achieved data fidelity, residual dirty images associated with zoom region (i) are shown in Fig. 9. The hyperspectral algorithms deliver noise-like residuals, enabling improved data fidelity over monochromatic algorithms. Unlike HyperAIRI and WSClean, the signature of the “dancing ghosts” is somewhat visible in the residual dirty image of Hyper-uSARA, suggesting a minor data over-fitting. The source’s spectral index maps are shown in Fig. 10. Hyperspectral algorithms yield smaller and more physically plausible spectral index values, while uSARA and AIRI show moderately higher values likely due to spectral inconsistency across channels. This is exemplified by the filaments associated with the two AGN and in their associated lobes.
@@
7 Conclusions
In this paper, we have introduced HyperAIRI, a novel PnP algorithm for extreme-scale hyperspectral RI imaging. HyperAIRI builds on the monochromatic FB-PnP algorithm AIRI, adapting its DNN denoisers to process hyperspectral RI image cubes, thereby exploiting their inherent spectral correlations. We also formally introduce the optimisation-based counterpart algorithm, Hyper-uSARA, which adopts the same FB iterative structure. Moreover, Hyper-uSARA can be interpreted as a simplified variant of HyperSARA (Abdulaziz et al., 2019; Thouvenin et al., 2023a), benefiting from an unconstrained formulation of the minimisation task, thus eliminating the need for exact prior knowledge of the RI noise level.
HyperAIRI denoisers take as input adjacent spectral channels to capture local correlations, while simultaneously injecting a power-law spectral model that facilitates the propagation of global spectral information across channels. This design reduces the need for global communications, thereby improving computational efficiency during both training and image reconstruction, while maintaining precision imaging. Furthermore, HyperAIRI denoisers adapt easily to any image cube size through its flexible spatial-spectral faceting. Inheriting from AIRI its training strategy, which is fully decoupled from the measurement process, HyperAIRI denoisers can be applied under varying observational setup without the need for retraining. HyperAIRI also benefits from convergence guarantees owing to its non-expensive denoisers. The resulting precision, flexibility and scalability mark a significant improvement over the state-of-the-art methods, so far struggling to achieve these requirements simultaneously.
@@ 7 Conclusions
Despite its promise for precision hyperspectral imaging, HyperAIRI remains computationally expensive due to the highly iterative nature of the underpinning FB, in which frequent calls of the measurement operator dominate the overall runtime. Inspired by CLEAN-based algorithms, exploring a major-minor cycle algorithmic structure with PSF approximations of the measurement operators inside minor cycles is a potential solution to reduce the computational cost. Additionally, building on the initial attempts in small-scale cases (Aghabiglou et al., 2025), developing a fully GPU-based implementation tailored for extreme-scale problems can significantly bridge the computational gap with WSClean.
Beyond scalability, further gains in reconstruction quality could be achieved. In particular, exploring advanced denoiser architectures that incorporate self- and cross-attention mechanisms (Vaswani et al., 2017), to exploit non-local features across the spatial and spectral domains, could enhance HyperAIRI’s precision. Endowing HyperAIRI with a calibration functionality is also crucial to ensure the high-precision imaging required for high-sensitivity observations from next-generation telescopes.
Acknowledgements
@@
Data Availability
HyperAIRI DNNs are available on the Heriot-Watt University research portal. The images used to generate training datasets are sourced as follows. Optical astronomy images are gathered from NOIRLab/NSF/AURA/H.Schweiker/WIYN/T.A.Rector (University of Alaska Anchorage). Medical images are obtained from the NYU fastMRI Initiative database (Zbontar et al., 2018; Knoll et al., 2020). The RI images used to build the simulated test dataset come from NRAO Archives, LOFAR HBA Virgo cluster survey (Edler et al., 2023), and LoTSS-DR2 survey (Shimwell et al., 2022). The ASKAP data underlying this article (calibrated visibilities and mosaic images of Scheduling Blocks) are made publicly available for viewing and downloading at CSIRO ASKAP Science Data Archive (CASDA, Chapman et al., 2015), and can be accessed with the unique Project Identifiers AS101. The uSARA and AIRI reconstructions of ASKAP used for comparison can be found through the DOI 10.17861/5a60f25b-d43b-4807-ba02-a69bc460db03.
@@ Appendix A Notation list
$\lambda$ The regularisation parameter in the objective function
@@ Appendix B WSClean Imaging Command
wsclean -join-channels -no-mf-weighting -fit-beam \
  -fit-spectral-pol 5 -channel-range 0 36 \
  -channels-out 36 -multiscale \
  -scale \${PIXELSIZE}asec -niter 6000000 \
  -nmiter 30 -weight briggs 0.0 -mgain 0.8 \
  -size 512 512 -auto-threshold 0.4 \
  -minuvw-m 0.001 -auto-mask 1 -mem 95 \${MSPATH}.
@@
  -size 4096 4096 -scale 2.2asec -minuvw-m 60  \
  -no-mf-weighting -fit-spectral-pol 4 \
  -channel-range 0 288 -channels-out 8 -mem 95 \
  -weight briggs -0.25 -weighting-rank-filter 3 \
@@
  -no-min-grid-resolution -auto-threshold 1.0 \
  -auto-mask 2.5 -mgain 0.8 -fit-beam -pol i \

### v2

@@ VI.1 Data details for the field SB9442-35
Figure 6: Reconstructed images of the field SB9442-35 at the frequencies 817 MHz (a, c) and 1069 MHz (b, d), obtained with AIRI (a, b), and uSARA (c, d) (A. G. Wilber et al., 2023b, a). Each panel is overlaid with zooms on selected regions: zoom (i) is centred on the “dancing ghosts”, zooms (ii) and (iv) show regions with both extended and compact radio sources, zoom (iii) contains the star-forming galaxy NGC 7090, zooms (v)-(viii) focus on four compact sources.
@@
Figure 7: Reconstructed images of the field SB9442-35 at the frequencies 817 MHz (a) and 1069 MHz (b), obtained using WSClean with joint-channel deconvolution. The zoom regions are identical to those in Figure 6.
The hyperspectral measurements used in this experiment are part of the ASKAP EMU Pilot Survey (R. P. Norris et al., 2021). In this survey, each Scheduling Block (SB) provides 36 different measurement sets, each corresponding to a different primary beam. For the real data experiment, we select the measurement set of Beam 35 of SB9442, centered at the radio source PKS 2130-538 of the galaxy cluster Abell 3785, known as the “dancing ghosts”. The observation spans 10 hours covering the frequency band 800-1088 MHz with a frequency step size of 1 MHz, yielding 288 spectral channels. The selected hyperspectral measurements were calibrated with the ASKAPsoft (A. Hotan et al., 2021) and are accessible through the CSIRO ASKAP Science Data Archive (J. Chapman et al., 2015).
VI.2 Imaging settings
Following the imaging considerations of A. G. Wilber et al. (2023a, b), the measurement set is divided into 8 spectral windows, combining 36 channels each. The visibilities in each spectral window are concatenated together during imaging. The target image cube thus contains 8 spectral channels. Each channel has a pixel size of $4096\times 4096$ and a cell size of 2.20 arcsec, achieving a final field of view of $2.57^{\circ}\times 2.57^{\circ}$, corresponding to a super-resolution factor of 2.22 at the highest-frequency channel and 2.91 at the lowest-frequency channel. The $w$-correction is enabled for all algorithms to account for $w$-effect in wide-field imaging. Briggs weighting with the robustness parameter fixed to 0.25-0.25 is considered.
The choices of algorithm-specific parameters are detailed below. For WSClean, we enable the joint-channel deconvolution and consider 4 polynomial terms for spectral fitting. The $w$-gridder is selected for gridding and $w$-correction. The remaining parameters are set consistent with the considerations of A. G. Wilber et al. (2023a), and the full WSClean command is shown in Appendix B. For comparison, the WSClean image cube is normalized channel-wise by the flux of the corresponding restoring beam as per (24). The regularization parameter $\lambda$ in Hyper-uSARA and the effective noise level for the denoisers in HyperAIRI are adjusted to 0.3 and 0.5 times the heuristic noise level (14) respectively. This adjustment improves the overall imaging quality while mitigating calibration artifacts. The MATLAB-based measurement operator encapsulates a hybrid $w$-projection $w$-stacking approach to correct for the $w$-effect (A. Dabbech et al., 2022, 2025).
To evaluate the spectral index maps from the reconstructed image cubes, we follow the considerations of A. G. Wilber et al. (2023a, b). For each image cube, a smoothing Gaussian kernel was applied, with a full width at half maximum (FWHM) of 20 arcsec for WSClean and 5 arcsec for the remaining algorithms. Subsequently, the logarithm of the pixel intensities of these channels are fitted to a linear function with the slope to be the spectral index map, as shown in (23).
@@ VI.3 Reconstruction results
Figure 8: Reconstructed images of the field SB9442-35 at the frequencies 817 MHz (a, c) and 1069 MHz (b, d), obtained with HyperAIRI (a, b) and Hyper-uSARA (c, d). The zoom regions are identical to those in Figure 6.
@@
We examine the reconstructions provided by the different algorithms in Figure 6-8, at both lowest (panels (a) and (c)) and highest (panels (b) and (d)) frequencies. Figure 6 exhibits AIRI (panels (a) and (b)) and uSARA (panels (c) and (d)) images from A. G. Wilber et al. (2023b, a). Figure 7 presents WSClean images. Figure 8 displays HyperAIRI (panels (a) and (b)) and Hyper-uSARA (panels (b) and (d)) images. In each panel, we consider zooms on eight selected regions as follows: region (i) centered at the “dancing ghosts”; regions (ii) and (iv) corresponding to areas containing both extended and compact radio sources; region (iii) centered on the star-forming galaxy NGC 7090; and regions (v)-(viii) centered on compact sources with varying morphologies.
@@
Compared with their monochromatic counterparts, the proposed hyperspectral algorithms recover significantly more faint background sources, capture finer details of the radio emission, and exhibit improved inter-channel consistency. Among the hyperspectral approaches, HyperAIRI achieves distinguished source sharpness and recovers the largest number of compact sources, albeit with some noise-like residual patterns suggesting some extent of data over-fitting. On the other hand, Hyper-uSARA fails to capture some faint compact background sources due to its joint sparsity regularization, enforcing pixel values of sources, detected in some channels only, to be set to zero. WSClean reconstructions are smoother by design, with some faint sources appear to be buried in the noise. While one could further optimize the noise-level scaling factors on a per-channel basis as in A. G. Wilber et al. (2023a, b) or consider per-channel visibility calibration to better balance background residuals against source recovery, these processing strategies lie beyond the scope of this experiment.
@@
A detailed examination of the zoom regions highlights substantial differences in the imaging precision of the different algorithms in terms of angular resolution, dynamic range and the ability to capture faint emission, and spectral correlation. Firstly, the superior angular resolution achieved by HyperAIRI at both channels compared to the benchmark is exemplified by zoom regions (vi), (vii) and (viii), with Hyper-uSARA ranking second. Both algorithms reveal inner cores and jets of the selected radio galaxies, whereas their monochromatic counterparts and even WSClean provide a blurred depiction of the same sources. Notably, HyperAIRI resolves the inner core of the radio galaxy in region (vi), and distinctly separates the two compact sources at the center of region (viii). Secondly, the enhanced sensitivity and dynamic-range of the proposed hyperspectral algorithms is showcased in zoom region (v). Both HyperAIRI and Hyper-uSARA successfully recover the faint jets emanating from the two hotspots. The faint features are completely absent in the monochromatic reconstructions, and are not revealed clearly in WSClean’s images. Thirdly, consistency of spectral behavior achieved by the proposed hyperspectral algorithms is exemplified by the V-shaped emission in region (ii), the star-forming galaxy NGC 7090 in region (iii), and the two-sided jet emission and nearby point sources in region (iv), with faint structures preserved across channels. Examining region (iii) at the highest frequency, monochromatic algorithms fail to recover NGC 7090’s fusiform morphology. For WSClean, the source is buried in the noise. By contrast, both HyperAIRI and Hyper-uSARA reliably restore NGC 7090’s elongated form, with the former providing a clearer depiction of the diffuse emission.
@@
Figure 9: Residual dirty images zoomed at the “dancing ghosts”, for channel 1 (top row) and 8 (bottom row) from the reconstructions produced by the algorithms shown in Figure 6-8
@@
Figure 10: The spectral index maps of the “dancing ghosts” evaluated from the reconstructions of different algorithms. Panels (a) and (b) are extracted from the reconstructions of A. G. Wilber et al. (2023a, b). The maps are computed after convolving channel images with a Gaussian beam of 20 arcsec for WSClean and 5 arcsec for the other reconstructions.
In what follows, we provide a deeper analysis of the recovery of the “dancing ghosts” in zoom region (i), including data fidelity and the source’s spectral index map. The source is the most prominent in this field, notable for its complex morphology. It comprises the emission of two active galactic nuclei (AGN): the first (WISEA J213406.70-533418.7) at the center of the north bridge, and the second (WISEA J213417.69-533811.1) in the southeast filament (R. P. Norris et al., 2021; V. Velović et al., 2023). All hyperspectral algorithms provide consistent reconstructions across channels. In particular, two jets from the two lobes of the first AGN respectively pointing towards the northeast exhibit similar shapes in both the first and last channels. Hyper-uSARA offers sharper edges and lower background noise, albeit with minor wavelets artifacts around the jets and lobes of the second AGN. HyperAIRI tends to deliver slightly higher resolution and recover more of the neighboring compact point-like sources than Hyper-uSARA. However, the PnP algorithm seems to depict the north bridge with a chain of dots, which is likely a reconstruction artifact. This could be indicative of training dataset bias, with half of the images simulated from optical astronomical images dominated by point-like sources. For insight into the achieved data fidelity, residual dirty images associated with zoom region (i) are shown in Figure 9. The hyperspectral algorithms deliver noise-like residuals, enabling improved data fidelity over monochromatic algorithms. Unlike HyperAIRI and WSClean, the signature of the “dancing ghosts” is somewhat visible in the residual dirty image of Hyper-uSARA, suggesting a minor data over-fitting. The source’s spectral index maps are shown in Figure 10. Hyperspectral algorithms yield smaller and more physically plausible spectral index values, while uSARA and AIRI show moderately higher values likely due to spectral inconsistency across channels. This is exemplified by the filaments associated with the two AGN and in their associated lobes.
@@
VII Conclusions
In this paper, we have introduced HyperAIRI, a novel PnP algorithm for extreme-scale hyperspectral RI imaging. HyperAIRI builds on the monochromatic FB-PnP algorithm AIRI, adapting its DNN denoisers to process hyperspectral RI image cubes, thereby exploiting their inherent spectral correlations. We also formally introduce the optimization-based counterpart algorithm, Hyper-uSARA, which adopts the same FB iterative structure. Moreover, Hyper-uSARA can be interpreted as a simplified variant of HyperSARA (A. Abdulaziz et al., 2019; P.-A. Thouvenin et al., 2023a), benefiting from an unconstrained formulation of the minimization task, thus eliminating the need for exact prior knowledge of the RI noise level.
HyperAIRI denoisers take as input adjacent spectral channels to capture local correlations, while simultaneously injecting a power-law spectral model that facilitates the propagation of global spectral information across channels. This design reduces the need for global communications, thereby improving computational efficiency during both training and image reconstruction, while maintaining precision imaging. Furthermore, HyperAIRI denoisers adapt easily to any image cube size through its flexible spatial-spectral faceting. Inheriting from AIRI its training strategy, which is fully decoupled from the measurement process, HyperAIRI denoisers can be applied under varying observational setup without the need for retraining. HyperAIRI also benefits from convergence guarantees owing to its non-expansive denoisers. The resulting precision, flexibility and scalability mark a significant improvement over the state-of-the-art methods, so far struggling to achieve these requirements simultaneously.
@@ VII Conclusions
Despite its promise for precision hyperspectral imaging, HyperAIRI remains computationally expensive due to the highly iterative nature of the underpinning FB, in which frequent calls of the measurement operator dominate the overall runtime. Inspired by CLEAN-based algorithms, exploring a major-minor cycle algorithmic structure with PSF approximations of the measurement operators inside minor cycles is a potential solution to reduce the computational cost. Additionally, building on the initial attempts in small-scale cases (A. Aghabiglou et al., 2025), developing a fully GPU-based implementation tailored for extreme-scale problems can significantly bridge the computational gap with WSClean.
Beyond scalability, further gains in reconstruction quality could be achieved. In particular, exploring advanced denoiser architectures that incorporate self- and cross-attention mechanisms (A. Vaswani et al., 2017), to exploit non-local features across the spatial and spectral domains, could enhance HyperAIRI’s precision. Endowing HyperAIRI with a calibration functionality is also crucial to ensure the high-precision imaging required for high-sensitivity observations from next-generation telescopes.
Data Availability
HyperAIRI DNNs are available on the Heriot-Watt University research portal. The images used to generate training datasets are sourced as follows. Optical astronomy images are gathered from NOIRLab/NSF/AURA/H.Schweiker/WIYN/T.A.Rector (University of Alaska Anchorage). Medical images are obtained from the NYU fastMRI Initiative database (J. Zbontar et al., 2018; F. Knoll et al., 2020). The RI images used to build the simulated test dataset come from NRAO Archives, LOFAR HBA Virgo cluster survey (H. Edler et al., 2023), and LoTSS-DR2 survey (T. Shimwell et al., 2022). The ASKAP data underlying this article (calibrated visibilities and mosaic images of Scheduling Blocks) are made publicly available for viewing and downloading at CSIRO ASKAP Science Data Archive (CASDA, J. Chapman et al., 2015), and can be accessed with the unique Project Identifiers AS101. The uSARA and AIRI reconstructions of ASKAP used for comparison can be found through the DOI 10.17861/5a60f25b-d43b-4807-ba02-a69bc460db03.
@@ Appendix A Notation list
$\lambda$ The regularization parameter in the objective function
@@ Appendix B WSClean Imaging Command
wsclean -join-channels -no-mf-weighting \
  -fit-beam -fit-spectral-pol 5 \
  -channel-range 0 36 -channels-out 36 \
  -multiscale -scale \${PIXELSIZE}asec \
  -niter 6000000 -nmiter 30 -weight \
  briggs 0.0 -mgain 0.8 -size 512 512 \
  -auto-threshold 0.4 -minuvw-m 0.001 \
  -auto-mask 1 -mem 95 \${MSPATH}.
@@
  -size 4096 4096 -scale 2.2asec \
  -minuvw-m 60 -no-mf-weighting \
  -fit-spectral-pol 4 -channel-range 0 288 \
  -channels-out 8 -mem 95 \
  -weight briggs -0.25 \
  -weighting-rank-filter 3 \
@@
  -no-min-grid-resolution \
  -auto-threshold 1.0 -auto-mask 2.5 \
  -mgain 0.8 -fit-beam -pol i \